# Sharp error bounds for edge-element discretisations of the high-frequency Maxwell equations

### v2

@@ Appendix A The Maxwell radial PML problem
This section recaps the definition of the Maxwell radial PML problem from [19], [56, §13.5.3.2, Page 378], [4] (using slightly different notation) and shows that the coeffcients $\mu$ and $\epsilon$ in this case satisfy (2.15) (see Lemma A.1 below).
@@
and $E_{\rm scat}$ satisfies the Silver-Müller radiation condition (see, e.g., [56, Equation 1.29]).
@@
Let $R_{{\rm tr}}>R_{\rm PML,-}>R_{\rm scat}$ and let $\Omega_{{\rm tr}}\subset\mathbb{R}^{d}$ be a bounded Lipschitz open set with $B_{R_{{\rm tr}}}\subset\Omega_{{\rm tr}}\subset B_{CR_{{\rm tr}}}$ for some $C>0$ (i.e., $\Omega_{{\rm tr}}$ has characteristic length scale $R_{{\rm tr}}$). Let $\Omega:=\Omega_{{\rm tr}}\cap\Omega_{+}$. For $20\leq\theta<\pi/$ 2, let the PML scaling function $f_{\theta}\in C^{1}([0,\infty);\mathbb{R})$ be defined by $f_{\theta}(r):=f(r)\tan\theta$ for some $f$ satisfying
@@
i.e., the scaling “turns on” at $r=R_{\rm PML,-}$, and is linear when $r\geq R_{\rm PML,+}$. Note that $R_{{\rm tr}}$ can be $<R_{\rm PML,+}$, i.e., truncation can occur before linear scaling is reached. Given $f_{\theta}(r)$, let
@@
$\mu:=\begin{cases}\mu_{\rm scat}&\text{ in }B_{R_{\rm PML,-}},\\ HDH^{T}&\text{ in }(B_{R_{\rm PML,-}})^{c}\end{cases}\text{ and }\epsilon:=\begin{cases}\epsilon_{\rm scat}&\text{ in }B_{R_{\rm PML,-}},\\ HDH^{T}&\text{ in }(B_{R_{\rm PML,-}})^{c}\end{cases}$ (A.3)
@@
We denote by $\widehat{I}^{\rm c},\widehat{I}^{\rm d}$ the canonical Nédélec and Raviart-Thomas interpolants of degree $p$ on the reference element $\widehat{K}$ (see, e.g., [56, §5.4–5.5], [23, Chapter 16]). We then consider the interpolation operators $I^{\rm c}_{h}:Z^{2}({\cal T}_{h})\to V_{h}$ and $I^{\rm d}_{h}:H^{1}_{\rm pw}(\mathcal{T}_{h})\to V_{h}$ by setting
@@
$\|\widehat{v}-\widehat{I}^{\rm c}\widehat{v}\|_{L^{2}(\widehat{K})}\leq C\left(|\widehat{v}|_{H^{\ell}(\widehat{K})}+|{\rm curl}\,\widehat{v}|_{H^{\ell}(\widehat{K})}\right).$
@@
$\|v-I^{\rm c}_{h}v\|_{L^{2}(K)}\leq Ch_{K}^{-1}h_{K}^{3/2}\left(|\widehat{v}|_{H^{\ell}(\widehat{K})}+|{\rm curl}\,\widehat{v}|_{H^{\ell}(\widehat{K})}\right).$ (B.10)
@@
Since $I^{\rm d}$ is continuous over $H^{1}(\widehat{K})$ and preserves polynomials of degree $p-1$, the Bramble-Hilbert lemma (see, e.g., [18, Theorem 28.1], [23, §11.3]) implies that there exists $C>0$ such that, for $\ell\in\{1,\ldots,p\}$,

### v3

@@ Appendix A The Maxwell radial PML problem
This section recaps the definition of the Maxwell radial PML problem from [19], [57, §13.5.3.2, Page 378], [4] (using slightly different notation) and shows that the coeffcients $\mu$ and $\epsilon$ in this case satisfy (2.15) (see Lemma A.1 below).
@@
and $E_{\rm scat}$ satisfies the Silver-Müller radiation condition (see, e.g., [57, Equation 1.29]).
@@
Let $R_{{\rm tr}}>R_{\rm PML,-}>R_{\rm scat}$ and let $\Omega_{{\rm tr}}\subset\mathbb{R}^{d}$ be a bounded Lipschitz open set with $B_{R_{{\rm tr}}}\subset\Omega_{{\rm tr}}\subset B_{CR_{{\rm tr}}}$ for some $C>0$ (i.e., $\Omega_{{\rm tr}}$ has characteristic length scale $R_{{\rm tr}}$). Let $\Omega:=\Omega_{{\rm tr}}\cap\Omega_{+}$. For $0\leq\theta<\pi/2$, let the PML scaling function $f_{\theta}\in C^{1}([0,\infty);\mathbb{R})$ be defined by $f_{\theta}(r):=f(r)\tan\theta$ for some $f$ satisfying
@@
i.e., the scaling “turns on” at $r=R_{\rm PML,-}$, and is linear when $r\geq R_{\rm PML,+}$. Note that $R_{{\rm tr}}$ can be $R_{\rm PML,+}$, i.e., truncation can occur before linear scaling is reached. Given $f_{\theta}(r)$, let
@@
[…]
@@
We denote by $\widehat{I}^{\rm c},\widehat{I}^{\rm d}$ the canonical Nédélec and Raviart-Thomas interpolants of degree $p$ on the reference element $\widehat{K}$ (see, e.g., [57, §5.4–5.5], [23, Chapter 16]). We then consider the interpolation operators $I^{\rm c}_{h}:Z^{2}({\cal T}_{h})\to V_{h}$ and $I^{\rm d}_{h}:H^{1}_{\rm pw}(\mathcal{T}_{h})\to V_{h}$ by setting
@@
[…]
@@
[…]
@@
[…]